# Towards Full-scene Domain Generalization in Multi-agent Collaborative Bird’s Eye View Segmentation for Connected and Autonomous Driving

### v2

@@ III-B Amplitude Augmentation
$x^{s\rightarrow t}_{i}=\mathbf{F}^{-1}\left(A(x_{j}^{s\rightarrow t}),P(x^{s}_{i})\right)$ (2)
Discussion. Here we elaborate the rationale behind the proposed method. Several factors contribute to this choice, with the first being its independence of annotation. In those scenarios related to CAVs, acquiring labeled data is both costly and labor-intensive, often cited as a bottleneck in research and development, and posing challenges for both industry and academia, especially when accounting for time-varying weather conditions (e.g., sunny, rainy, foggy, snowy), color discrepancies, and diverse driving scenarios. Within the domain of collaborative BEV segmentation, open source datasets are scarce. To our best knowledge, only two datasets exist xuOPV2VOpenBenchmark2022 ; liV2XSimMultiAgentCollaborative2022 , both of which are derived from simulations rather than real-world environments. The lack of real-world data underscores the urge need for developing methods that can operate effectively in the absence of extensive annotated datasets. Based on this observation and restriction, we propose AmpAug without using additional annotation. The second notable advantage of our proposed method is its plug-and-use nature. This feature is particularly advantageous as it allows for effortless integration into existing training pipelines without any modifications. This plug-and-use characteristic not only enhances the usability of the method but also significantly reduces the time and resources required for implementation.
@@ III-C Meta-consistency Training
Meta training. We use the gradient-based meta-learning algorithm as our foundational learning scheme, which can learn generalizable model by simulating the real-world domain shift in the training phase. In our settings, the domain shift comes from the data generated from the frequency space with amplitude augmentation. Specifically, in each iteration, we consider the raw camera data perceived by CAVs’ cameras as meta training data, and its counterparts generated from frequency space as meta testing data. The meta-learning can be divided into two stages. We first update the model parameters $\theta$ using the meta-train data $\mathcal{D}_{s}$ with cross entropy $\mathcal{L}_{CE}$:
$\mathcal{L}_{CE}=\frac{1}{N}\sum^{N-1}_{i=0}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right)$ (3)
@@
$\theta^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{CE}(\mathcal{D}_{s};\theta)$ (4)
where $\alpha$ is the learning rate in the inner loop update. Then, we leverage a meta-learning step to enhance the generalization ability of the model by simulating the domain shift in the meta-test data $\mathcal{D}_{t}$. Specifically, in the meta-learning phase, the meta-objective $\mathcal{L}_{meta}$ is computed with the updated parameters $\theta^{\prime}$, but optimized towards the original parameters $\theta$. Intuitively, apart from learning the segmentation task on the meta-train data $\mathcal{D}_{s}$, this learning scheme teach the model to learn how to adapt to the domain shift across meta-train data $\mathcal{D}_{s}$ and meta-test data $\mathcal{D}_{t}$.
Consistency Optimization. While the meta-training scheme can simulate domain shifts and help the model adaptation to these shifts, it primarily focuses on making the model robust under shifts within the source domain (as simulated with meta-source and meta-target domains). However, this approach alone is not sufficient to ensure that the model learn the representations that are invariant under different domains. When training a model generalizable for unseen domains, it is crucial to guide the model in learning representations that remain consistent across multiple domains. This is important because, in the feature space, all the labeled data from source domains contribute to training a model that can generalize to previously unseen domains liDomainGeneralizationAdversarial2018 . If the data distributions across various domains within a given feature space remain dissimilar, it hinders the model’s ability to effectively generalize to unseen domains. This is why we propose a consistency optimization method to force the model to learn the domain-invariant representations.
@@
$\hat{y}_{i}^{s}=f(x_{i}^{s};\theta)=D_{seg}(E(x_{i}^{s};\theta_{E});\theta_{D})$ (5)
where model parameters $\theta$ are decoupled into $\theta_{E}$ and $\theta_{D}$ for encoder and decoder, respectively. Then, after computing the meta-train loss $\mathcal{L}_{CE}(x_{i}^{s};\theta)$, we can obtain the updated model parameters $\theta^{\prime}$ by the inner loop update (Here, the updated model parameters $\theta^{\prime}$ will just be used in the meta-test step, and will be released after current iteration, the real updated parameters will be computed later). Then we leverage the updated model parameters $\theta^{\prime}$ to compute the meta-test forward propagation prediction on target domain images $x^{t}_{i}$ which are generated by our amplitude augmentation method:
$\hat{y}_{i}^{t}=f(x_{i}^{t};\theta^{\prime})=D_{seg}(E(x_{i}^{t};\theta_{E}^{\prime});\theta_{D}^{\prime})$ (6)
Then, we extract the latent representation from input images $x_{i}^{s}$ and target domain images $x_{i}^{t}$ in meta-train and meta-test step, respectively:
$z^{s}=E(x_{i}^{s};\theta_{E}),~{}z^{t}=E(x_{i}^{t};\theta_{E}^{\prime})$ (7)
Next, in order to enhance the domain-invariant feature capture capability of the model, we design the consistency loss function that leverages the maximum mean discrepancy grettonKernelMethodTwoSampleProblem to make the latent representations $z^{s}$ and $z^{t}$ more similar. We first map the features to a reproducing kernel Hilbert space (RKHS) $\mathcal{F}$, and then compute the distance between their mean in RKHS as follows:
$\mathcal{L}_{cons}=\left\|\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}\phi(z_{i}^{s})-\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\phi(z_{i}^{t})\right\|^{2}_{\mathcal{F}}$ (8)
where $\phi$ is the kernel function that maps the distribution to RKHS, $n_{s}$ and $n_{t}$ are the numbers of samples in source and target domain, respectively. Typically, the Gaussian RBF kernel is used as the kernel function:
$k(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right)$ (9)
Then, the consistency loss function is given by grettonKernelMethodTwoSampleProblem :
$\begin{split}\mathcal{L}_{cons}&=\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}k(z_{i}^{s},z_{j}^{s})+\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}k(z_{i}^{t},z_{j}^{t})\\ &-\frac{2}{n_{s}n_{t}}\sum_{i,j=1}^{n_{s},n_{t}}k(z_{i}^{s},z_{j}^{t})\end{split}$ (10)
@@
Overall learning objective. The overall meta learning objective is composed of the cross entropy loss $\mathcal{L}_{CE}$ and the consistency loss $\mathcal{L}_{cons}$ as follows:
$\mathcal{L}_{meta}=\mathcal{L}_{CE}(x_{i}^{t};\theta^{\prime})+\beta\mathcal{L}_{cons}(x_{i}^{s},x_{i}^{t};\theta^{\prime})$ (11)
where $\beta$ is the balancing weight on the consistency loss, while $\theta^{\prime}$ is the updated model parameters obtained from the inner loop update in Eq. (4). Then, the inner loop objective and the meta learning objective are optimized together with respect to the original parameters $\theta$ as follows:
$\hat{\theta}=\theta-\gamma\nabla\left(\mathcal{L}_{CE}(x_{i}^{s};\theta)+\mathcal{L}_{meta}(x_{i}^{s},x_{i}^{t};\theta^{\prime})\right)$ (12)
@@ III-D Intra-system Domain Alignment During Inference
In collaborative perception for CAV driving, we observe that the ego vehicle and other vehicles are situated at different environments, e.g., unbalanced lighting conditions: one vehicle in the shade and the other in the open. Moreover, different types of car cameras can cause chromatic aberration. These phenomena can result in intra-system domain discrepancy, which can degrade the performance of a collaborative perception system, to some extent. In this subsection, we propose a simple yet effective method to align the intra-system domain shift during inference. This method can translate the image style by unifying the pixel distribution in color space. Inspired by reinhardColorTransferImages2001 , we employ image translation based on the LAB color space (LAB color space is another color model) to minimize domain discrepancy. Similar to RGB, LAB color space is also a three-dimensional color space. The three dimensions of LAB color space are $L^{*}$, $a^{*}$ and $b^{*}$. $L^{*}$ represents the brightness of the color, and $a^{*}$ and $b^{*}$ represent the color components. This approach is motivated by two key observations. Firstly, the LAB color space has a broader gamut compared to the RGB color space. Secondly, as demonstrated in Fig. 3, images translated in the LAB color space more resemble the style of target domain images than those translated directly in the RGB color space. Thus, after considering a source domain RGB image $x^{s}$ and a target domain RGB image $x^{t}$, we convert the RGB image to the XYZ color space by:
$\left[\begin{array}[]{c}X\\ Y\\ Z\end{array}\right]=\left[\begin{array}[]{lll}0.4124&0.3575&0.1804\\ 0.2126&0.7151&0.0721\\ 0.0193&0.1191&0.9502\end{array}\right]\left[\begin{array}[]{l}R\\ G\\ B\end{array}\right]$ (13)
@@
$L^{\star}$ $=116f\left(Y/Y_{n}\right)-16$ (14)
$a^{\star}$ $=500\left[f\left(X/X_{n}\right)-f\left(Y/Y_{n}\right)\right]$
$b^{\star}$ $=200\left[f\left(Y/Y_{n}\right)-f\left(Z/Z_{n}\right)\right]$
where $f(t)=t^{1/3}$, if $t>\left(\frac{6}{29}\right)^{3}$, otherwise $f(t)=\frac{1}{3}\left(\frac{29}{6}\right)^{2}t+\frac{4}{29}$, and $X_{n},Y_{n},Z_{n}$ are the coordinate values of the reference white point. The LAB color space is a perceptually uniform color space, which means that a small change in LAB color space corresponds to a small change in the perceived color. Therefore, we can translate the image style by manipulating the LAB color space. After converting the RGB images to the LAB color space, denoted as $x^{s}_{LAB}$ and $x^{t}_{LAB}$, respectively, we first compute the mean and standard deviation of the LAB color channels of $x^{s}_{LAB}$ and $x^{t}_{LAB}$, denoted as $\mu^{s}_{LAB}$, $\sigma^{s}_{LAB}$, $\mu^{t}_{LAB}$, $\sigma^{t}_{LAB}$, respectively. Then, we can obtain the translated image $x^{s}_{LAB}\rightarrow x^{t}_{LAB}$ by:
$\hat{x}^{s}_{LAB}=\sigma^{t}_{LAB}/\sigma^{s}_{LAB}\cdot(x^{s}_{LAB}-\mu^{s}_{LAB})+\mu^{t}_{LAB}$ (15)
@@
In summary, in ollaborative perception, the ego vehicle first converts its perceived RGB images to LAB color space and obtains $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$. Then, the ego vehicle shares $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$ with other vehicles. Other vehicles can then use these parameters to translate their perceived RGB images to the target domain style. This process helps reduce the domain discrepancy among the collaborative vehicles. After that, other vehicles extract the intermediate features from the translated images and send these features to the ego vehicle. The ego vehicle then fuses these features to predict the final segmentation mask.
Figure 4: Visualization of the foggy dataset. The first row is the original image, the second row is the synthesized foggy image.
@@
Figure 7: Visualization of the night dataset, the first raw and the third raw is the original images, the second raw and the fourth raw is the synthesized night images.
@@ IV Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed method. Specifically, we choose OPV2V as our fundamental dataset and synthesize four different datasets with different domain discrepancy. We first introduce the datasets and metrics used in our experiments. Then, we compare our method with several state-of-the-art methods. Finally, we conduct ablation studies to evaluate the effectiveness of our method.

### v3

@@ III-B Amplitude Augmentation
$x^{s\rightarrow t}_{i}=\mathbf{F}^{-1}\left(A(x_{j}^{s\rightarrow t}),P(x^{s}_{i})\right).$ (3)
Discussion. Here we elaborate the rationale behind the proposed method. Several factors contribute to this design, with the first being its independence of annotation. In those scenarios related to CAVs, acquiring labeled data is both costly and labor-intensive, often cited as a bottleneck in research and development, and posing challenges for both industry and academia, especially when accounting for time-varying weather conditions (e.g., sunny, rainy, foggy, snowy), color discrepancies, and diverse driving scenarios. Within the domain of collaborative BEV segmentation, open-source datasets are scarce. To our best knowledge, only two datasets exist xuOPV2VOpenBenchmark2022 ; liV2XSimMultiAgentCollaborative2022 , both of which are derived from simulators rather than real-world environments. The lack of real-world data underscores the urge need to develop methods that can operate effectively in the absence of extensive annotated datasets. Based on this observation and restriction, we propose AmpAug without using additional annotation. The second notable advantage of our proposed method is its plug-and-use nature. This feature is particularly advantageous as it allows for effortless integration into existing training pipelines without anymodifications. This plug-and-use characteristic not only enhances the usability of the method but also significantly reduces the time and resources required for implementation.
In addition, here we also elaborate on the choice of $\alpha$ in Eq. (2) and the impact of selecting different proportions $\alpha$. As for the mask $M$, we have $M(h,w)=\mathbf{1}_{h\times w}$, where $h\in[-\alpha H:\alpha H]$, $w\in[-\alpha W:\alpha W]$, where $\alpha\in(0,1)$ controls the size of the mask. From Eq. (2), we can see that the mask $M$ controls the proportion of the source and target images. When $\alpha=0$, the image $x^{s\rightarrow t}$ remains identical to the original source image $x^{s}$. On the other hand, when $\alpha=1.0$, the amplitude of $x^{s}$ is completely substituted with the amplitude of $x^{t}$. We have observed in our practice that as $\alpha$ progresses towards 1.0, the image $x^{s\rightarrow t}$ increasingly resembles the target image $x^{t}$. However, this transition also introduces noticeable artifacts in the image. Therefore, we set $\alpha=0.01$ to balance the trade-off between the source and target images.
@@ III-C Meta-consistency Training
Meta-training. We use the gradient-based meta-learning algorithm as our foundational learning scheme, which can learn a generalizable model by simulating the real-world domain shift in the training phase. In our settings, the domain shift comes from the data generated from the frequency domain with amplitude augmentation. Specifically, in each iteration, we consider the raw camera source data $\mathcal{D}_{s}$ perceived by CAVs’ cameras as meta-training data and its counterparts generated from frequency domain by AmpAug as meta-testing data $\mathcal{D}_{s\rightarrow t}$. The meta-learning can be divided into two stages. We first update the model parameters $\theta$ using the meta-training data $\mathcal{D}_{s}$ with cross entropy $\mathcal{L}_{CE}$:
$\mathcal{L}_{CE}=\frac{1}{N}\sum^{N-1}_{i=0}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right),$ (4)
@@
$\theta^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{CE}(\mathcal{D}_{s};\theta),$ (5)
where $\alpha$ is the learning rate in the inner loop update. Then, we leverage a meta-learning step to enhance the generalization ability of the model by simulating the domain shift in the meta-test data $\mathcal{D}_{t}$. Specifically, in the meta-learning phase, the meta-objective $\mathcal{L}_{meta}$ is computed with the updated parameters $\theta^{\prime}$, but optimized towards the original parameters $\theta$. Intuitively, apart from learning the segmentation task on the meta-train data $\mathcal{D}_{s}$, this learning scheme teaches the model to learn how to adapt to the domain shift across meta-training data $\mathcal{D}_{s}$ and meta-testing data $\mathcal{D}_{s\rightarrow t}$.
Consistency Optimization. While the meta-training scheme can simulate domain shifts and help the model adapt to these shifts, it primarily focuses on making the model robust under shifts within the source domain (as simulated with meta-source and meta-target domains). However, this approach alone is not sufficient to ensure that the model learns the representations that are invariant under different domains. When training a model generalizable for unseen domains, it is crucial to guide the model in learning representations that remain consistent across multiple domains. This is important because, in the feature space, all the labeled data from source domains contribute to training a model that can generalize to previously unseen domains liDomainGeneralizationAdversarial2018 . If the data distributions across various domains within a given feature space remain dissimilar, it hinders the model’s ability to effectively generalize to unseen domains. This is why we propose a consistency optimization method to force the model to learn the domain-invariant representations.
@@
$\hat{y}_{i}^{s}=f(x_{i}^{s};\theta)=D_{seg}(E(x_{i}^{s};\theta_{E});\theta_{D}),$ (6)
where model parameters $\theta$ are decoupled into $\theta_{E}$ and $\theta_{D}$ for encoder and decoder, respectively. Then, after computing the meta-training loss $\mathcal{L}_{CE}(x_{i}^{s};\theta)$, we can obtain the updated model parameters $\theta^{\prime}$ by the inner loop update (Here, the updated model parameters $\theta^{\prime}$ will just be used in the meta-testing step, and will be released after current iteration, the real updated parameters will be computed later). Then, we leverage the updated model parameters $\theta^{\prime}$ to compute the meta-testing forward propagation prediction on target domain images $x^{t}_{i}\in\mathcal{D}_{s\rightarrow t}$ that are generated by our amplitude augmentation method:
$\hat{y}_{i}^{t}=f(x_{i}^{t};\theta^{\prime})=D_{seg}(E(x_{i}^{t};\theta_{E}^{\prime});\theta_{D}^{\prime}).$ (7)
Next, we extract the latent representation from input images $x_{i}^{s}$ and target domain images $x_{i}^{t}$ in the meta-training and meta-testing steps, respectively:
$z^{s}=E(x_{i}^{s};\theta_{E}),~{}z^{t}=E(x_{i}^{t};\theta_{E}^{\prime}).$ (8)
In order to enhance the domain-invariant feature capture capability of the model, we design the consistency loss function that leverages the maximum mean discrepancy grettonKernelMethodTwoSampleProblem2006 to make the latent representations $z^{s}$ and $z^{t}$ more similar. We first map the features to a reproducing kernel Hilbert space (RKHS) $\mathcal{F}$, and then compute the distance between their mean in RKHS as follows:
$\mathcal{L}_{cons}=\left\|\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}\phi(z_{i}^{s})-\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\phi(z_{i}^{t})\right\|^{2}_{\mathcal{F}},$ (9)
where $\phi$ is the kernel function that maps the distribution to RKHS, and $n_{s}$ and $n_{t}$ are the numbers of samples in the source and target domain, respectively. Typically, the Gaussian RBF kernel is used as the kernel function:
$k(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right).$ (10)
Then, the consistency loss function is given by grettonKernelMethodTwoSampleProblem2006 :
$\begin{split}\mathcal{L}_{cons}&=\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}k(z_{i}^{s},z_{j}^{s})+\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}k(z_{i}^{t},z_{j}^{t})\\ &-\frac{2}{n_{s}n_{t}}\sum_{i,j=1}^{n_{s},n_{t}}k(z_{i}^{s},z_{j}^{t}).\end{split}$ (11)
@@
Overall learning objective. The overall meta-learning objective is composed of the cross entropy loss $\mathcal{L}_{CE}$ and the consistency loss $\mathcal{L}_{cons}$ as follows:
$\mathcal{L}_{meta}=\mathcal{L}_{CE}(x_{i}^{t};\theta^{\prime})+\beta\mathcal{L}_{cons}(x_{i}^{s},x_{i}^{t};\theta^{\prime}),$ (12)
where $\beta$ is the balancing weight on the consistency loss, while $\theta^{\prime}$ are the updated model parameter obtained from the inner loop update in Eq. (5). Then, the inner loop objective and the meta-learning objective are optimized together with respect to the original parameters $\theta$ as follows:
$\hat{\theta}=\theta-\gamma\nabla\left(\mathcal{L}_{CE}(x_{i}^{s};\theta)+\mathcal{L}_{meta}(x_{i}^{s},x_{i}^{t};\theta^{\prime})\right).$ (13)
@@ III-D Intra-system Domain Alignment During Inference
In collaborative perception for CAV driving, we observe that the ego CAV and other CAVs are situated at different environments, e.g., unbalanced lighting conditions: one CAV in the shade and the other in the open. Moreover, different types of car cameras can cause chromatic aberration. These phenomena can result in intra-system domain discrepancy, which can degrade the performance of a collaborative perception system, to some extent. In this subsection, we propose a simple yet effective method to align the intra-system domain shift during inference. This method can translate the image style by unifying the pixel distribution in color space. Inspired by reinhardColorTransferImages2001 , we employ image translation based on the LAB color space (LAB color space is another color model) to minimize domain discrepancy. Similar to RGB, LAB color space is also a three-dimensional color space. The three dimensions of the LAB color space are $L^{*}$, $a^{*}$, and $b^{*}$, respectively, where $L^{*}$ represents the brightness of the color, and $a^{*}$ and $b^{*}$ represent the color components. This approach is motivated by two key observations. Firstly, the LAB color space has a broader gamut compared to the RGB color space. Secondly, as demonstrated in Fig. 4, images translated in the LAB color space more resemble the style of target domain images than those translated directly in the RGB color space. Thus, after considering a source domain RGB image $x^{s}$ and a target domain RGB image $x^{t}$, we convert the RGB image to the XYZ color space by:
$\left[\begin{array}[]{c}X\\ Y\\ Z\end{array}\right]=\left[\begin{array}[]{lll}0.4124&0.3575&0.1804\\ 0.2126&0.7151&0.0721\\ 0.0193&0.1191&0.9502\end{array}\right]\left[\begin{array}[]{l}R\\ G\\ B\end{array}\right].$ (14)
@@
$L^{\star}$ $=116f\left(Y/Y_{n}\right)-16,$ (15)
$a^{\star}$ $=500\left[f\left(X/X_{n}\right)-f\left(Y/Y_{n}\right)\right],$
$b^{\star}$ $=200\left[f\left(Y/Y_{n}\right)-f\left(Z/Z_{n}\right)\right],$
where $f(t)=t^{1/3}$, if $t>\left(\frac{6}{29}\right)^{3}$, otherwise $f(t)=\frac{1}{3}\left(\frac{29}{6}\right)^{2}t+\frac{4}{29}$, while $X_{n},Y_{n},Z_{n}$ are the coordinate values of the reference white point. The LAB color space is a perceptually uniform color space, which means that a small change in the LAB color space corresponds to a small change in the perceived color. Therefore, we can translate the image style by modifying the numerical values in the LAB color space. After converting the RGB images to the LAB color space, denoted as $x^{s}_{LAB}$ and $x^{t}_{LAB}$, respectively, we first compute the mean and standard deviation of the LAB color channels of $x^{s}_{LAB}$ and $x^{t}_{LAB}$, denoted as $\mu^{s}_{LAB}$, $\sigma^{s}_{LAB}$, $\mu^{t}_{LAB}$, $\sigma^{t}_{LAB}$, respectively. Then, we can obtain the translated image $x^{s}_{LAB}\rightarrow x^{t}_{LAB}$ by:
$\hat{x}^{s}_{LAB}=\sigma^{t}_{LAB}/\sigma^{s}_{LAB}\cdot(x^{s}_{LAB}-\mu^{s}_{LAB})+\mu^{t}_{LAB}.$ (16)
@@
In summary, in collaborative perception, the ego CAV first converts its perceived RGB images to the LAB color space and obtains $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$. Then, the ego CAV shares $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$ with other CAVs. Other CAVs can then use these parameters to translate their perceived RGB images to the target domain style. This process helps reduce the domain discrepancy among the collaborative CAVs. After that, other CAVs extract the intermediate features from the translated images and send these features to the ego vehicle. The ego CAV then fuses these features to predict the final segmentation mask.
Figure 5: Visualization of the foggy dataset. The first row is the original image, and the second row is the synthesized foggy image.
@@
Figure 8: Visualization of the night dataset, the first row and the third row are the original images while the second row and the fourth row are the synthesized night images.
@@ IV Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed method. Specifically, we choose OPV2V as our fundamental dataset and synthesize four different datasets with different domain discrepancies. We first introduce the datasets and metrics used in our experiments in Sec. IV-A. Then, we present the implementation details of our method in Sec. IV-B. Next, we compare our method with several state-of-the-art methods in Sec. IV-C. Finally, we conduct ablation studies to evaluate the effectiveness of our method in Sec. IV-D.